OpenStudy (anonymous):

HELP MEDAL!!! What is the 3rd term when -2h4 + 5h2 - 3 - 2h - 8h3 is arranged in descending order?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

i like your picture. The scion FRS is a torquey car

OpenStudy (anonymous):

yeah man i have one now but my dream is to Rocket Bunny mine

OpenStudy (anonymous):

any help on this question?

OpenStudy (anonymous):

@haydenhd

OpenStudy (anonymous):

so its descending which means going down.

OpenStudy (anonymous):

i think H is the third term

OpenStudy (anonymous):

alright thanks man!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

could you help me with a graphing question @haydenhd

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is the equation of the graph below? y = - (x - 3)2 + 1 y = - (x + 1)2 + 4 y = (x + 1)2 - 4 y = (x - 4)2 + 1 http://gyazo.com/9d3920f59040e5cec4bdeeef0964f4e5

OpenStudy (anonymous):

Well this is an application of pythagorean's theorem so we need to find the slope.

OpenStudy (anonymous):

I think the slope is 2

OpenStudy (anonymous):

I think A.

OpenStudy (anonymous):

Any reasoning?

OpenStudy (anonymous):

cause for every one block the y axis moves, the x axis moves three. And the only answer that has a three in it is A.

OpenStudy (anonymous):

What grade are you in

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

Given f(x) = the quantity of x minus 7, divided by 3, solve for f-1(4). -1 19 33 37 Help?

OpenStudy (anonymous):

And you have a car?

OpenStudy (anonymous):

Yeah im 16

OpenStudy (anonymous):

Oh, so you have your permit.

OpenStudy (anonymous):

thats cool

OpenStudy (anonymous):

License

OpenStudy (anonymous):

i am sixteen and saving up for a jeep

OpenStudy (anonymous):

i want to raise it, get bigger tires, then spread out the tires.

OpenStudy (anonymous):

Yeah 4 inch lifts nice body kit on

OpenStudy (anonymous):

wrangler right?

OpenStudy (anonymous):

Yup... Do you need any more help

OpenStudy (anonymous):

Given f(x) = the quantity of x minus 7, divided by 3, solve for f-1(4). -1 19 33 37

OpenStudy (anonymous):

I dont know this one,, sorry

OpenStudy (anonymous):

ok how about this http://gyazo.com/9ae380ab43e3643af6ef458979069141

OpenStudy (anonymous):

Definately D

OpenStudy (anonymous):

There is a lot of math to it that would take to long to explain

OpenStudy (anonymous):

just the answer then please?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

ohh i see thanks man